What happens to takt time if demand rate is doubled?

  • Cannot be determined
  • Increases by 50%
  • Decreases by 50%
  • No change
Answer #1

Takt Time = 1/Demand rate

If Demand rate is X, then takt time = 1/x

Now if demand rate becomes 2X, takt time = 1/2X = 0.5/X

So Takt time reduces by 0.5/X / 1/x = 50%